OpenStudy (ayeeeitskateee):

URGENT: Need help badly!! Will medal/fan A sphere has a volume of 4500π cm3. What is the length of its radius?

Vocaloid (vocaloid):

well, volume of a sphere is 4/3*pi*r^3, right? we can set the given volume 4500*pi = 4/3*pi*r^3 what's left after you divide both sides by (4/3)*pi?

OpenStudy (ayeeeitskateee):

375? @Vocaloid

Vocaloid (vocaloid):

not quite, check your math again

OpenStudy (ayeeeitskateee):

Well, how do you do it? I took 4/3 and divided it by 4500.

OpenStudy (ayeeeitskateee):

@Vocaloid

Vocaloid (vocaloid):

4500/(4/3) = ?

Vocaloid (vocaloid):

plug it into your calculator if you aren't sure

OpenStudy (ayeeeitskateee):

This time it says 3375.